# Joint multi-dimensional dynamic attention and transformer for general image restoration

### v1

@@ IV-G Discussions
| Method | Venue | Denoising (CBSD68 [64]) | Dehazing | Deraining | Average | FLOPs (G)/ #Params (M) | ||
| SOTS-Outdoor [57] | Rain100L [18] | |||||||
| MPRNet [28] | 21’CVPR | 33.27/0.920 | 30.76/0.871 | 27.29/0.761 | 28.00/0.958 | 33.86/0.958 | 30.63/0.894 | 148.55/3.64 |
| Restormer [10] | 22’CVPR | 33.72/0.930 | 30.67/0.865 | 27.63/0.792 | 27.78/0.958 | 33.78/0.958 | 30.75/0.901 | 140.99/26.13 |
| NAFNet [46] | 22’ECCV | 33.03/0.918 | 30.47/0.865 | 27.12/0.754 | 24.11/0.928 | 33.64/0.956 | 29.67/0.844 | 63.33/67.89 |
| MambaIR [73] | 24’ECCV | 33.88/0.931 | 30.95/0.874 | 27.74/0.793 | 29.57/0.970 | 35.42/0.969 | 31.51/0.907 | 229.80/26.78 |
| MDDA-former | - | 33.94/0.933 | 31.28/0.888 | 28.05/0.798 | 29.17/0.968 | 35.26/0.969 | 31.54/0.911 | 67.38/25.92 |
[…]
@@ IV-H Ablation Study
| Model | chunk in MDAB | chunk in ETB | FLOPs (G) | #Params (M) | PSNR (rain) |
| (a) | ✗ | ✗ | 43.15 | 17.29 | 37.56 |
| (b) | ✓ | ✗ | 42.32 | 16.76 | 37.42 |
| (c) | ✗ | ✓ | 40.86 | 16.63 | 37.33 |

### v2

@@ IV-G Discussions
Method | Venue | Denoising (CBSD68 [70]) | Dehazing | Deraining | Average |
SOTS-Outdoor [63] | Rain100L [18] | |||||||||
MPRNet [30] | 21’CVPR | 33.27/0.920 | 30.76/0.871 | 27.29/0.761 | 28.00/0.958 | 33.86/0.958 | 30.63/0.894 | 148.55/3.64 | ||
Restormer [10] | 22’CVPR | 33.72/0.930 | 30.67/0.865 | 27.63/0.792 | 27.78/0.958 | 33.78/0.958 | 30.75/0.901 | 140.99/26.13 | ||
NAFNet [52] | 22’ECCV | 33.03/0.918 | 30.47/0.865 | 27.12/0.754 | 24.11/0.928 | 33.64/0.956 | 29.67/0.844 | 63.33/67.89 | ||
MambaIR [79] | 24’ECCV | 33.88/0.931 | 30.95/0.874 | 27.74/0.793 | 29.57/0.970 | 35.42/0.969 | 31.51/0.907 | 229.80/26.78 | ||
MDDA-former | - | 33.94/0.933 | 31.28/0.888 | 28.05/0.798 | 29.17/0.968 | 35.26/0.969 | 31.54/0.911 | 67.38/25.92 | ||
AirNet [35] | 22’CVPR | 33.92/0.933 | 31.26/0.888 | 28.00/0.797 | 27.94/0.962 | 34.90/0.967 | 31.20/0.910 | 301.27/5.77 | ||
@@ IV-H Ablation Study
Model |
FLOPs (G) | #Params (M) | PSNR (rain) | ||||
(a) | ✗ | ✗ | 43.15 | 17.29 | 37.56 | ||||
(b) | ✓ | ✗ | 42.32 | 16.76 | 37.42 | ||||
(c) | ✗ | ✓ | 40.86 | 16.63 | 37.33 | ||||